John begins painting a house at 12:00pm. At 2:00 pm Ted joins John. They finish together at 6:00pm. It would taken John 9 hours to paint the house alone. How long would it have taken Ted to paint the house alone?
Answer by richwmiller(17219)

John begins painting a house at 12:00 noon. At 2:00 pm Ted joins John. They finish together at 6:00 pm. It would taken John 9 hours to paint the house alone. How long would it have taken Ted to paint the house alone?
6/9+4/t=1
4/t=1/3
ted=12 hours alone